OpenStudy (anonymous):

If a line with a slope of 4 crosses the y-axis at (0, 1), what is the equation of the line?

OpenStudy (anonymous):

90

OpenStudy (anonymous):

@dhiraj123

OpenStudy (anonymous):

?

OpenStudy (anonymous):

a.2x − y = -4 b.2x + y = -4 c.2x + y = 4 d.2x − y = 4

OpenStudy (anonymous):

what's the question

OpenStudy (anonymous):

If a line with a slope of 4 crosses the y-axis at (0, 1), what is the equation of the line?

OpenStudy (anonymous):

....a or c maybe

OpenStudy (anonymous):

Slope intercept of a line is \(y=mx+b\), where m is the slope, and b is the y-intercept. Based on the information that you have, the equation of the line will be \(y=4x+1\) Now, rearrange this equation into standard form \(Ax+By=C\). can you so this part?

OpenStudy (anonymous):

noo could you do it for me?

OpenStudy (anonymous):

Transpose \(4x\) on the left side or subtract both sides by \(4x\). BUT there is something wrong. are you sure you typed the question correctly? because if you'll change it to standard form, you'll get \(y-4x=1\) which is not in the choices.

OpenStudy (anonymous):

its 5

OpenStudy (anonymous):

oh im about to be bumped in 2:16